OpenStudy (anonymous):

what are the solutions of 3x^2-4=0?

OpenStudy (phi):

if the equation were 3x - 4 =0 could you solve for x?

OpenStudy (anonymous):

4/3

OpenStudy (phi):

you do the same thing, except you get x^2 = 4/3

OpenStudy (phi):

now take the square root of both sides

OpenStudy (anonymous):

is it x=2√3

OpenStudy (phi):

you start with 4/3 so you get ±sqrt(4/3) or ±2/sqrt(3) (divide by sqrt(3) ) however, people don't like sqrt in the denominator, so they multiply top and bottom by sqrt of 3 like this \[ \frac{2}{\sqrt{3}}\cdot \frac{\sqrt{3}}{\sqrt{3}}= \frac{2\sqrt{3}}{3}\] and it is plus or minus (both will work in the equation) \[ ± \frac{2\sqrt{3}}{3}\]

OpenStudy (anonymous):

oh ok i see

OpenStudy (phi):

in case it is not obvious \[ \sqrt{\frac{4}{3}} \text{ can be written as }\frac{\sqrt{4}}{\sqrt{3}}= \frac{2}{\sqrt{3}}\]

OpenStudy (anonymous):

ok. thanks a lot for your help (: